OpenStudy (anonymous):

Television sets. What does it mean to refer to a 20 -in tv set or a 25-in tv set? Such units refer to the diagonal of the screen. a 35-in tv set also has a width of 28 inches. What is the height of a 35-inch tv ...inches???

OpenStudy (amistre64):

h = sqrt(35^2 - 28^2)

OpenStudy (amistre64):

h = 21

OpenStudy (anonymous):

\[h=\sqrt{35^2-28^2}\] \[h=\sqrt{1225-784}\] \[h=\sqrt{441}\] \[h=21\]

OpenStudy (anonymous):

damn i am slow

OpenStudy (anonymous):

at any rate i hope it is clear that this all comes from \[a^2+b^2=c^2\] where a and b are the sides and c is the hypotenuse. so you are thinking of the diagonal as the hypotenuse and the sides as 35 and 28

OpenStudy (anonymous):

actually i meant you are thinking of the hypotenuse as 35 and one of the sides at 28

OpenStudy (anonymous):

so you get \[28^2+h^2=35^2\] and then solve for h

OpenStudy (anonymous):

\[h^2=35^2-28^2\] etc

OpenStudy (anonymous):

Ok i see what you are saying

OpenStudy (anonymous):

good!

OpenStudy (anonymous):

if one leg of a triangle is 5 and the hypotenuse is 13 then you know \[5^2+b^2=13^2\]

OpenStudy (anonymous):

solving for b gives first \[b^2=13^2-5^2\] and then \[b=\sqrt{13^2-5^2}\]

OpenStudy (anonymous):

but if you do this often you just go right to the last step

OpenStudy (anonymous):

\[h=28 \text{Tan}\left[\text{ArcCos}\left[\frac{28}{35}\right]\right] \]